OpenStudy (anonymous):

2(1−k)x2 −4kx+k=0 Solve for x ?

OpenStudy (anonymous):

sorry, not written properly : 2(1−k)x^2 −4kx+k=0

OpenStudy (mimi_x3):

is this pertained to discriminants?

OpenStudy (anonymous):

Just to use quadratic formula, but don't know how to get it into ax^2+bx+c form c:

OpenStudy (anonymous):

It's in that form

OpenStudy (mimi_x3):

i think it's related to this thingy D = b^2 -4ac

OpenStudy (jhannybean):

\[2x^2 -2kx^2 -4kx+k=0\]\[x^2-kx^2 -2kx +\frac{k}{2}=0\]

OpenStudy (anonymous):

a = 2(1-k)

OpenStudy (jhannybean):

It is, mimi, but it had to be reduced first, lol

OpenStudy (jhannybean):

You can also use the completing the square method here.

OpenStudy (mimi_x3):

lel my bad...it has been almost 4 years since ive touched this

OpenStudy (jhannybean):

Mmhmm.

OpenStudy (anonymous):

I'm sorry, I still don't get it! What exactly is a, b, and c and how did you get it ?

OpenStudy (jhannybean):

Vertex form of a quadratic: \[y=\color{red}ax^2+\color{red}bx+\color{red}c\]your equation:\[y=\color{red}{2(1-k)}x^2\color{red}{-4k}x+\color{red}k=0\] Now just match up a,b,c from the quadratic to the variables in your equation, and then apply the quadratic formula, \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]Guess that would be the easiest way of going about it.

OpenStudy (jhannybean):

sorry, it's not the vertex form of the quadratic, forgot to erase that. It's just the parabolic form of the quadratic.